OpenStudy (anonymous):

Solutions To Equation? I Was In The Hospital When We Went Over This In Class, Can Someone Answer & Explain? 3(4x - 3) - 7x = 5x - 9 A.) None B.) One C.) Infinite Medal To Best Answer! :)!

OpenStudy (x3_drummerchick):

i believe it's infinite. 12x - 9 - 7x = 5x - 9 5x - 9 = 5x - 9 no matter what you put in there for x, it's going to equal the same on both sies of the equation

OpenStudy (x3_drummerchick):

substitute 0, for example, for x: 3( (4*0) - 3) - 7 (0) = 5 (0) - 9 3 (-3) - 0 = 0 - 9 -9 = -9

OpenStudy (x3_drummerchick):

one more example. substitute 1 for x: 3 ((4*1) - 3))- 7(1) = 5 (1) - 9 3 (4-3) - 7 = 5 - 9 3 (1) - 7 = 5 - 9 -4 = - 4

OpenStudy (anonymous):

Thank You Very Much! I Love Your Explanation!